# The Shape of Money Laundering: Subgraph Representation Learning on the Blockchain with the Elliptic2 Dataset

### v2

@@ Abstract.
ccs: Security and privacy Database activity monitoringccs: Computing methodologies Machine learningccs: Applied computing Network forensics
@@ 2.2.2. Graph construction
    Define the time window, the maximum number of hops to be traversed, and the early stopping condition when a change of ownership is likely to happen.
@@
    Annotate subgraphs for the final output, keeping only nodes that are part of either licit (licit subgraphs) or suspicious (suspicious subgraphs) paths.
The time window was chosen to be 1 year of blockchain data, and the maximum number of hops was 6. The transactions at step (3) were capped to a maximum for each cluster, to help balance the dataset for different actors with varying blockchain activity.
@@ 5.1. Insights from a cryptocurrency exchange service
According to the exchange, less than 0.1% of customer accounts have been associated with money laundering or fraud, whereas at least 26.9% of the accounts highlighted by the model predictions were found to have such an association. This suggests that our model performs very well.
@@ 6. Subgraph learning at scale
Subgraph methods like GLASS require only a few changes to those for node classification. Here, we introduce a scalable GNN training system, SALIENT (Kaler et al., 2022), together with its improved successor SALIENT++ (Kaler et al., 2023), which was designed for node classification workloads. Then, we discuss how systems like SALIENT/SALIENT++ can be adapted to address subgraph classification workloads.
@@ 6.2. Adapting node-classification GNNs to subgraph classification workloads
For subgraph methods such as GLASS, existing efficient training systems for node classification can be modified with ease to handle subgraph classification workloads. We take SALIENT and SALIENT++ for example.
@@ 7. Summary
We have contributed a second large, labeled cryptocurrency transaction data set, Elliptic2, five years after Elliptic1 was published. Elliptic2 is more than two orders of magnitude larger than Elliptic1 and it supports a different machine-learning task – subgraph classification – for AML. It offers research opportunities in not only financial forensics but also machine learning, particular, subgraph representation learning. We demonstrated its utility through training subgraph-based predictive models and discovered an effective one: GLASS.

### v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: KDD MLF ’24: KDD Workshop on Machine Learning in Finance; August 26, 2024; Barcelona, Spainbooktitle: KDD MLF ’24: KDD Workshop on Machine Learning in Finance, August 26, 2024, Barcelona, Spainccs: Security and privacy Database activity monitoringccs: Computing methodologies Machine learningccs: Applied computing Network forensics
@@ 2.2.2. Graph construction
    Define the time window, the maximum number of hops to be traversed, and an early stopping condition when a change of ownership is likely to happen (e.g., when an unknown cluster with high activity is found during traversal).
@@
    Annotate subgraphs for the final output, keeping only unknown nodes that are part of either licit (licit subgraphs) or suspicious (suspicious subgraphs) paths.
The time window was chosen to be 1 year of blockchain data, and the maximum number of hops was 6 for computational reasons. The transactions at step (3) were capped to a maximum for each cluster, to help balance the dataset for different actors with varying blockchain activity.
@@ 5.1. Insights from a cryptocurrency exchange service
According to the exchange, less than 0.1% of customer accounts have been associated with money laundering or fraud, whereas at least 26.9% of the accounts highlighted by the model predictions were found to have such an association. This suggests that the empirical precision of our model significantly surpasses that of a naïve random model.
@@ 6. Subgraph learning at scale
Subgraph methods like GLASS require some changes to those for node classification. Here, we introduce a scalable GNN training system, SALIENT (Kaler et al., 2022), together with its improved successor SALIENT++ (Kaler et al., 2023), which was designed for node classification workloads. Then, we discuss how systems like SALIENT/SALIENT++ can be adapted to address subgraph classification workloads.
@@ 6.2. Adapting node-classification GNNs to subgraph classification workloads
For subgraph methods such as GLASS, existing efficient training systems for node classification can be modified to handle subgraph classification workloads. We take SALIENT and SALIENT++ for example.
@@ 7. Summary
We have contributed a second large, labeled cryptocurrency transaction data set, Elliptic2, five years after Elliptic1 was published. Elliptic2 is more than two orders of magnitude larger than Elliptic1 and it supports a different machine-learning task – subgraph classification – for AML. It offers research opportunities in not only financial forensics but also machine learning, particularly, subgraph representation learning. We demonstrated its utility through training subgraph-based predictive models and discovered an effective one: GLASS.